OpenStudy (anonymous):

Can someone explain how to do this problem. I already know the answer, I just want to learn. 1. Solve 5x2 = –45x.

OpenStudy (anonymous):

5*2 =10 10=-45x divide both sides by -45, to isolate x x=-2/9

OpenStudy (anonymous):

since this is a quadratic equation, solve it like a quadratic equation... get it in standard form, then factor, then use the zero product property to find the solution(s).

OpenStudy (anonymous):

is it 5*2 or 5^2?

OpenStudy (anonymous):

it is 5x^2

OpenStudy (anonymous):

\[5x ^{2}\]

OpenStudy (anonymous):

5x^2=-45x +45x on both sides to balance 5x^2+45x=0 quadratic formula ithink...

OpenStudy (anonymous):

what is the answer?

OpenStudy (anonymous):

x=-9

OpenStudy (lgbasallote):

5x^2 + 45x = 0 divide both sides by 5 x^2 + 9x = 0 factor out x x(x+9) = 0 equate each factor to zero x = 0 AND x + 9 = 0 solve for x

OpenStudy (anonymous):

ohh yeah, forgot to factor out 5

OpenStudy (lgbasallote):

@destinyshaw there are two answers to this :)

OpenStudy (lgbasallote):

because it's quadratic

OpenStudy (anonymous):

the problem is : \[5x ^{2} = -45x\] that is the problem, not it equals 0

OpenStudy (anonymous):

wouldn't you use the quadratic formula?

OpenStudy (anonymous):

hey, its uj1024, not ujwall, thats a different user

OpenStudy (lgbasallote):

look at the above replies i just continued from there

OpenStudy (lgbasallote):

yeah i deleted lol :P i have bad eyesight

OpenStudy (anonymous):

.....I just need to know how to work out this problem :/

OpenStudy (lgbasallote):

quadratic formula is also applicable..both will show same results

OpenStudy (anonymous):

Ok, I just tired to work it out.. and it did not work, please someone explain to me in depth